A mass of 5KG is on top of a building of heights 10M calculate its potential energy the mass then falls at a speed of 8M/S in 4 seconds find its KE and momentum P what is the force and impulse the mass transfers to the ground
1
Expert's answer
2019-11-21T10:06:37-0500

Potential energy:


"E=mgh=(5)(9.8)(10)=490\ J"

"KE=0.5mv^2=0.5(5)(8)^2=160\ j"

"P=mv=(5)(8)=40\frac{kgm}{s}"

The impulse the mass transfers to the ground:


"I=P=40\frac{kgm}{s}"

The force:


"F=\frac{I}{t}=\frac{40}{4}=10\ N"